# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v4

@@ C.3 Canonical choice via a nondegenerate limit
$i_{x}:=-\frac{\beta_{x}}{2},\qquad i_{y}:=-\frac{\beta_{y}}{2}.$ (C.3)
These are exactly the choices obtained by eliminating the nilpotent component in (C.2).
C.4 Consistency check
With the canonical choice (C.3),
@@ C.5 Interpretation
The canonical choice (C.3) selects the derivatives that arise as limits of nondegenerate (elliptic or hyperbolic) structures and removes the nilpotent ambiguity. This provides a consistent notion of “constructive differentiability” in the parabolic setting.
Appendix D Jet Analysis of the $\varepsilon$–Family at the Constant Structure
This appendix analyzes the $\varepsilon$–family of rigid elliptic structures introduced in Chapter 4 from the viewpoint of jet theory at $\varepsilon=0$. Although the limiting structure is exactly the standard complex plane, the jet analysis reveals nontrivial infinitesimal geometry encoded in the family. The purpose of this appendix is to make precise what information is contained in the $\varepsilon$–jets and what conclusions can be drawn from them.
D.1 The $\varepsilon$–family and its constant limit
@@
D.2 First $\varepsilon$–jet of the coefficients
@@ D.3 Jet of the spectral parameter
D.4 Linearization of the Burgers transport
@@ D.5 Interpretation of the jet equation
D.6 What the jet analysis does not say
@@ D.7 Conceptual conclusion
D.8 Second $\varepsilon$–jet: onset of nonlinearity
@@
D.9 Third $\varepsilon$–jet: hierarchy of transport corrections
@@
D.10 Structure of the full jet hierarchy
@@
D.11 Refined conclusion of the jet analysis
@@
Appendix E Integrability, Coordinate Trivialization, and the Intrinsic Obstruction
@@
E.1 The Classical Integrability Objection
@@
E.2 What the Obstruction Does Not Measure
@@
E.3 Parallelism versus Coordinate Trivialization
@@
E.4 Eulerian and Lagrangian Viewpoints
@@ E.5 Conclusion
Appendix F Quaternionic Reformulation of Rigidity via the Hopf Fibration
F.1 Motivation
@@
$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0.$ (F.2)
@@
F.1.1 Normalization to a unit imaginary quaternion
Completing the square in (F.1) yields
@@ F.1.1 Normalization to a unit imaginary quaternion
$u(x,y):=\frac{2i(x,y)+\beta(x,y)}{\sqrt{\Delta(x,y)}}.$ (F.4)
@@
F.2 Hopf lift and gauge freedom
@@
$q\ \mapsto\ q\,e^{i_{0}\theta(x,y)}.$ (F.7)
F.3 Maurer–Cartan form and differentiation
@@
$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$ (F.8)
which lie in $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$. Differentiating (F.6) and using $(q^{-1})_{x}=-q^{-1}q_{x}q^{-1}$ gives
$u_{x}=[\Omega_{x},u],\qquad u_{y}=[\Omega_{y},u],$ (F.9)
@@
F.3.1 Vertical–horizontal decomposition
@@
$\Omega_{x}=\Omega_{x}^{\mathrm{ver}}+\Omega_{x}^{\mathrm{hor}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{ver}}+\Omega_{y}^{\mathrm{hor}},$ (F.10)
where $\Omega^{\mathrm{ver}}\parallel u$ and $\Omega^{\mathrm{hor}}\perp u$. Since $\Omega^{\mathrm{ver}}$ commutes with $u$, it follows from (F.9) that
$u_{x}=[\Omega_{x}^{\mathrm{hor}},u],\qquad u_{y}=[\Omega_{y}^{\mathrm{hor}},u].$ (F.11)
Under the gauge change (F.7), the Maurer–Cartan form changes by adding a vertical term, so the horizontal components and the commutators in (F.11) capture the gauge-invariant variation of $u$.
F.4 Quaternionic form of the obstruction
The obstruction $G=i_{x}+i\,i_{y}$ can be rewritten in terms of $u$. Up to the scalar normalizations in (F.4), the intrinsic combination governing deviations from the constant theory corresponds to
$\mathcal{G}:=u_{x}+u\,u_{y}.$ (F.12)
Using (F.11), we may express $\mathcal{G}$ in Hopf variables as
$\mathcal{G}=[\Omega_{x}^{\mathrm{hor}},u]\;+\;u\,[\Omega_{y}^{\mathrm{hor}},u].$ (F.13)
@@ F.4 Quaternionic form of the obstruction
F.5 Quaternionic rigidity theorem
Theorem F.1 (Quaternionic rigidity).
Let $u:\Omega\to S^{2}\subset\operatorname{Im}\mathbb{H}$ be the normalized generator (F.4) associated with an elliptic quadratic structure (F.1). Let $q$ be a local Hopf lift (F.6) with Maurer–Cartan components (F.8) and decomposition (F.10). Then the following are equivalent:
@@ F.5 Quaternionic rigidity theorem
    $u_{x}+u\,u_{y}=0\quad\text{on }\Omega.$ (F.14)
@@
Moreover, (F.14) is the Hopf/quaternionic avatar of the original rigidity condition $G=i_{x}+i\,i_{y}=0$.
@@
By (F.11), we have $u_{x}=[\Omega_{x}^{\mathrm{hor}},u]$ and $u_{y}=[\Omega_{y}^{\mathrm{hor}},u]$. Substituting these into (F.12) yields (F.13), and hence $\mathcal{G}\equiv 0$ is equivalent to (F.15). The gauge statement follows from the fact that changing the lift by (F.7) alters only the vertical part of the Maurer–Cartan form, leaving the horizontal commutator expressions unchanged. ∎
F.6 Interpretation
Equation (F.14) says that the $S^{2}$-valued imaginary direction field $u$ is transported compatibly with its own multiplication rule: along the characteristic direction $\partial_{x}+u\,\partial_{y}$, the field exhibits no genuine horizontal twisting in $S^{2}$, only Hopf fiber rotation. In this sense rigidity is a horizontality cancellation condition in the Hopf fibration rather than an analytic assumption.
Appendix G Rigidization as a Computational Normal Form
@@
G.1 Rigidization versus Uniformization
@@
G.2 Why “Weaker” Means “Easier”
@@
G.3 A Solver-Oriented Triage
@@ G.4 Rigidization as a Preconditioner
G.5 The Global Rigidization Problem
@@ G.7 Summary
Appendix H Symbolic Verification of the Rigid Second–Order Expansion
@@ 4 $d_{\bar z}$ f
Appendix I Symbolic Verification for the Rigidity-Flatness Theorem Condition

### v5

@@ B.3 Canonical choice via a nondegenerate limit
$i_{x}:=-\frac{\beta_{x}}{2},\qquad i_{y}:=-\frac{\beta_{y}}{2}.$ (B.3)
These are exactly the choices obtained by eliminating the nilpotent component in (B.2).
B.4 Consistency check
With the canonical choice (B.3),
@@ B.5 Interpretation
The canonical choice (B.3) selects the derivatives that arise as limits of nondegenerate (elliptic or hyperbolic) structures and removes the nilpotent ambiguity. This provides a consistent notion of “constructive differentiability” in the parabolic setting.
Appendix C Jet Analysis of the $\varepsilon$–Family at the Constant Structure
This appendix analyzes the $\varepsilon$–family of rigid elliptic structures introduced in Chapter 5 from the viewpoint of jet theory at $\varepsilon=0$. Although the limiting structure is exactly the standard complex plane, the jet analysis reveals nontrivial infinitesimal geometry encoded in the family. The purpose of this appendix is to make precise what information is contained in the $\varepsilon$–jets and what conclusions can be drawn from them.
C.1 The $\varepsilon$–family and its constant limit
@@
C.2 First $\varepsilon$–jet of the coefficients
@@ C.3 Jet of the spectral parameter
C.4 Linearization of the Burgers transport
@@ C.5 Interpretation of the jet equation
C.6 What the jet analysis does not say
@@ C.7 Conceptual conclusion
C.8 Second $\varepsilon$–jet: onset of nonlinearity
@@
C.9 Third $\varepsilon$–jet: hierarchy of transport corrections
@@
C.10 Structure of the full jet hierarchy
@@
C.11 Refined conclusion of the jet analysis
@@
Appendix D Integrability, Coordinate Trivialization, and the Intrinsic Obstruction
@@
D.1 The Classical Integrability Objection
@@
D.2 What the Obstruction Does Not Measure
@@
D.3 Parallelism versus Coordinate Trivialization
@@
D.4 Eulerian and Lagrangian Viewpoints
@@ D.5 Conclusion
Appendix E Quaternionic Reformulation of Rigidity via the Hopf Fibration
E.1 Motivation
@@
$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0.$ (E.2)
@@
E.1.1 Normalization to a unit imaginary quaternion
Completing the square in (E.1) yields
@@ E.1.1 Normalization to a unit imaginary quaternion
$u(x,y):=\frac{2i(x,y)+\beta(x,y)}{\sqrt{\Delta(x,y)}}.$ (E.4)
@@
E.2 Hopf lift and gauge freedom
@@
$q\ \mapsto\ q\,e^{i_{0}\theta(x,y)}.$ (E.7)
E.3 Maurer–Cartan form and differentiation
@@
$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$ (E.8)
which lie in $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$. Differentiating (E.6) and using $(q^{-1})_{x}=-q^{-1}q_{x}q^{-1}$ gives
$u_{x}=[\Omega_{x},u],\qquad u_{y}=[\Omega_{y},u],$ (E.9)
@@
E.3.1 Vertical–horizontal decomposition
@@
$\Omega_{x}=\Omega_{x}^{\mathrm{ver}}+\Omega_{x}^{\mathrm{hor}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{ver}}+\Omega_{y}^{\mathrm{hor}},$ (E.10)
where $\Omega^{\mathrm{ver}}\parallel u$ and $\Omega^{\mathrm{hor}}\perp u$. Since $\Omega^{\mathrm{ver}}$ commutes with $u$, it follows from (E.9) that
$u_{x}=[\Omega_{x}^{\mathrm{hor}},u],\qquad u_{y}=[\Omega_{y}^{\mathrm{hor}},u].$ (E.11)
Under the gauge change (E.7), the Maurer–Cartan form changes by adding a vertical term, so the horizontal components and the commutators in (E.11) capture the gauge-invariant variation of $u$.
E.4 Quaternionic form of the obstruction
The obstruction $G=i_{x}+i\,i_{y}$ can be rewritten in terms of $u$. Up to the scalar normalizations in (E.4), the intrinsic combination governing deviations from the constant theory corresponds to
$\mathcal{G}:=u_{x}+u\,u_{y}.$ (E.12)
Using (E.11), we may express $\mathcal{G}$ in Hopf variables as
$\mathcal{G}=[\Omega_{x}^{\mathrm{hor}},u]\;+\;u\,[\Omega_{y}^{\mathrm{hor}},u].$ (E.13)
@@ E.4 Quaternionic form of the obstruction
E.5 Quaternionic rigidity theorem
Theorem E.1 (Quaternionic rigidity).
Let $u:\Omega\to S^{2}\subset\operatorname{Im}\mathbb{H}$ be the normalized generator (E.4) associated with an elliptic quadratic structure (E.1). Let $q$ be a local Hopf lift (E.6) with Maurer–Cartan components (E.8) and decomposition (E.10). Then the following are equivalent:
@@ E.5 Quaternionic rigidity theorem
    $u_{x}+u\,u_{y}=0\quad\text{on }\Omega.$ (E.14)
@@
Moreover, (E.14) is the Hopf/quaternionic avatar of the original rigidity condition $G=i_{x}+i\,i_{y}=0$.
@@
By (E.11), we have $u_{x}=[\Omega_{x}^{\mathrm{hor}},u]$ and $u_{y}=[\Omega_{y}^{\mathrm{hor}},u]$. Substituting these into (E.12) yields (E.13), and hence $\mathcal{G}\equiv 0$ is equivalent to (E.15). The gauge statement follows from the fact that changing the lift by (E.7) alters only the vertical part of the Maurer–Cartan form, leaving the horizontal commutator expressions unchanged. ∎
E.6 Interpretation
Equation (E.14) says that the $S^{2}$-valued imaginary direction field $u$ is transported compatibly with its own multiplication rule: along the characteristic direction $\partial_{x}+u\,\partial_{y}$, the field exhibits no genuine horizontal twisting in $S^{2}$, only Hopf fiber rotation. In this sense rigidity is a horizontality cancellation condition in the Hopf fibration rather than an analytic assumption.
Appendix F Rigidization as a Computational Normal Form
@@
F.1 Rigidization versus Uniformization
@@
F.2 Why “Weaker” Means “Easier”
@@
F.3 A Solver-Oriented Triage
@@ F.4 Rigidization as a Preconditioner
F.5 The Global Rigidization Problem
@@ F.7 Summary
Appendix G Symbolic Verification of the Rigid Second–Order Expansion
@@ 4 $d_{\bar z}$ f
Appendix H Symbolic Verification for the Rigidity-Flatness Theorem Condition